I have two math related questions in need of explanation. 1rst:x:6x+8-3x+3=x-1. 2nd: After a 20%discount and 6% tax, a guitar cost $530.00. What was the original price?
Answer by rfer(16322)

A)
3x+11=x-1
3x-x=-1-11
2x=-12
x=-12/2
x=-6
B)
530/1.06=500
500/0.80=$625